# An Investigation of Robustness of LLMs in Mathematical Reasoning: Benchmarking with Mathematically-Equivalent Transformation of Advanced Mathematical Problems

### v1

@@ Introduction
5:   accept $v_{t}$ {two‑in‑a‑row passed}
6:   break
7:  else if $\mathbf{z}_{t}=\mathbf{1}$ then
8:   keep $v_{t}$ for next round
9:  else
10:   apply LLM‑suggested patch $\to v_{t}$
11:  end if
12:end for
13: human audit 10 % of accepted variants
Implementation Overview
Code release.
To facilitate double‑blind reviewing we publish only the subset of data (100 randomly chosen examples). An automated evaluator, putnam-cli.py, receives the names of target solver model and grader model and variant type to test. Supported back‑ends are (i) any HuggingFace‐compatible checkpoint via transformers, (ii) a local vllm server, or (iii) API clients including OpenAI, Gemini, Anthropic and OpenRouter Full data and generation scripts will be released post‑decision.
Surface generation.
Renaming variants are produced on a CPU‑only node by streaming o3 API calls. A five‑stage exponential‑back‑off retry (max 5 attempts, doubling timeout each time) masks transient API latency. Processing all 1 0511\,051 items in parallel takes $\sim$15 min wall‑clock.
Core‑step generation.
Kernel variant synthesis is more expensive because of multi‑turn chain‑of‑thought reasoning: end‑to‑end runtime is $\leq\!3\,\text{h}$ for the full corpus on a single 8‑core CPU, dominated by the 15‑iteration repair‑and‑verify loop.
PutnamGAP Dataset
Data Sources, Extraction & Annotation
Our benchmark comprises all Putnam Problems 1938–2024 (N=1 051N=1\,051 items after deduplication). The front-matter of every book contains the same fair-use clause, excerpted verbatim below:
“Individual readers … are permitted to make fair use of the material, such as to copy select pages for use in teaching or research.”
This clause grants us the legal right to reproduce problems and solutions for non-commercial academic evaluation. In line with AMS policy, we distribute only machine-readable IDs and LaTeX texts; raw PDF scans remain under the original AMS license, and any further redistribution must be cleared through the Copyright Clearance Center.
Original scans are processed via a 3-stage OCR routine: (i) Manual segmentation for every question-answer pair. (ii) MathPix for formula‐aware PDF-to-LaTeX conversion followed by (iii) custom post-filters that merge multi-line expressions and fix 4.2 % residual symbol errors. Each item is manually spot-checked ($\leq$2 min per problem) to ensure semantic fidelity before variant generation. Complete corpus list, OCR accuracy study, and cleaning scripts appear in Appendix D.
Dataset Statistics
Overall scale and balance.
Figure 2: Problem Topics and Classes
The benchmark comprises 1 051 original Putnam problems from 1938–2024 and five mathematically equivalent transformations, yielding 6 306 items. Part distribution is balanced (527 A vs. 524 B), and the canonical identifier $\langle\textit{year},\textit{part}\{A,B\},\textit{index}\rangle$ provides a difficulty proxy. Using indices 121\!\!-\!2 as Easy, 343\!\!-\!4 as Medium, and 565\!\!-\!6 as Hard, the corpus contains 32.3 % Easy, 32.3 % Medium, 32.2 % Hard, plus a 3.0 % extra–hard tail (indices 7–8).
Topic coverage.
Automatic tags in _meta.tag indicate broad mathematical coverage—Algebra (641), Analysis (521), Number Theory (392), Combinatorics (286), and Geometry (239).
Statement length.
After stripping  markup, an average statement is 35.7 tokens (min 0, max 152); all transformation families alter length by <2<\!2%, confirming that robustness drops cannot be explained by verbosity alone.
Quality control.
Every item has undergone single‐pass manual validation; inter‐rater Cohen $\kappa$ is not yet available and will be released in a future update.
Experimental Setup
Model Pool & Prompting
We evaluated 18 models (see 1 or Appendix A for a complete list).All models are queried under a unified zero‑shot template. A system instruction designates the model as “an expert mathematician” and asks it to show all work, while the user message embeds the problem. See Appendix for our full prompt. We fix temperature=0, top_p=1, and max_tokens=32000 or maximum token amount available in case some models have max_tokens maximum smaller than 32000. for every run except OpenAI O-series which require temperature=1. Solutions are then re‑submitted to a second template that grades the answer: a strict proof rubric for proof items and a lenient numeric rubric for calculation items. Both grader prompts require structured JSON output containing a binary grade field plus detailed feedback. Complete prompt code and hyper‑parameter settings are available in Appendix B and the anonymous repository (loader/prompt.py).
Scoring & Auto-Grader
We partition tasks into computational and proof-based categories and evaluate them with distinct graders.
Computational. Each candidate answer is normalized (whitespace, units, LaTeX macros) and passed to two scoring paths: (i) a strict string match against the reference solution; (ii) a latent grader—an LLM prompted to return ‘‘CORRECT’’ or ‘‘INCORRECT’’ given the reference answer and a rubric that disallows partial credit. We adopt path (ii) to mitigate formatting artifacts; if the two paths disagree we mark the item for manual audit (¡1 % of cases).
Proof-based. We provide the grader with an aligned, step-by-step reference proof and ask it to assign a binary grade plus a natural-language justification. Any skipped logical step or missing citation triggers a fail. A random 10 % sample is double-checked by independent volunteers; grader precision/recall is >>97 %.
All grader prompts, rubrics, and decision thresholds are released as supplementary material after the review period.
Compute & Reproducibility
All inference were performed through publicly available APIs. Each model was queried exactly once per item with the hyper-parameters in Table 2. Runs were executed from a single Ubuntu 22.04 host (11th Gen Intel(R) Core(TM) i7-11800H @ 2.30GHz); no local GPU was used. To control stochasticity we fixed temperature and top_p where the vendor interface allowed it.
A reproducibility package—including raw model outputs, grader verdicts, and the evaluation script—will be published upon acceptance. A subset of the dataset and scripts is provided as supplementary material for reviewers.
Results & Analysis
Table 1: Model Accuracy Rates across Categories (Percent Scale)
Model Original Descriptive ($\Delta$) Confusing ($\Delta$) Misleading ($\Delta$) Garbled String ($\Delta$) Kernel Variant ($\Delta$)
claude-opus-4 26.5∗∗∗ 23.0∗∗ (–3.5) 22.2∗∗ (–4.3) 21.7∗∗ (–4.8) 21.4∗∗ (–5.1) 13.8∗∗∗ (–12.7)
claude-sonnet-4 23.0∗∗∗ 20.6∗∗∗ (–2.5) 19.8∗∗ (–3.2) 18.6∗∗ (–4.4) 18.1∗∗∗ (–4.9) 11.1∗∗∗ (–11.9)
deepseek-prover 15.5∗∗∗ 15.2∗∗∗ (–0.3) 14.0∗∗∗ (–1.5) 12.8∗∗ (–2.7) 13.7∗∗∗ (–1.8) 9.2∗∗∗ (–6.3)
gemini-2.5-flash-lite 19.8∗∗∗ 18.8∗∗∗ (–0.9) 16.1∗∗ (–3.7) 15.8∗∗ (–4.0) 15.1∗∗∗ (–4.7) 6.6∗∗∗ (–13.2)
gemini-2.5-pro 78.4∗∗∗ 75.2∗∗ (–3.1) 74.3∗∗ (–4.1) 72.8∗∗∗ (–5.6) 72.9∗∗∗ (–5.4) 63.5∗∗∗ (–14.9)
gemini-2.5-flash 42.8∗∗∗ 42.6∗∗∗ (–0.2) 39.0∗∗ (–3.8) 40.9∗∗∗ (–1.9) 37.6∗∗ (–5.2) 27.6∗∗∗ (–15.2)
gpt-4.1 24.9∗∗∗ 23.4∗∗∗ (–1.5) 21.2∗∗ (–3.7) 22.0∗∗∗ (–2.9) 22.6∗∗∗ (–2.3) 14.8∗∗∗ (–10.1)
gpt-4.1-mini 28.6∗∗∗ 26.9∗∗∗ (–1.7) 27.1∗∗∗ (–1.5) 24.0∗∗ (–4.6) 25.1∗∗ (–3.4) 19.2∗∗∗ (–9.4)
gpt-4.1-nano 8.8∗∗∗ 8.9∗∗∗ (+0.1) 6.4∗∗ (–2.4) 7.3∗∗∗ (–1.5) 6.6∗∗∗ (–2.2) 6.8∗∗∗ (–2.0)
gpt-4o 6.5∗∗∗ 6.3∗∗∗ (–0.1) 5.3∗∗∗ (–1.1) 4.7∗∗ (–1.8) 4.7∗∗ (–1.8) 3.0∗∗∗ (–3.4)
gpt-4o-mini 4.3∗∗∗ 3.5∗∗∗ (–0.8) 2.5∗∗ (–1.8) 3.3∗∗∗ (–1.0) 3.2∗∗∗ (–1.1) 1.7∗∗∗ (–2.6)
grok4 60.1∗∗∗ 59.0∗∗∗ (–1.1) 56.6∗∗∗ (–3.4) 55.9∗∗ (–4.2) 55.2∗∗ (–4.8) 45.5∗∗∗ (–14.6)
kimi-k2 27.2∗∗∗ 25.8∗∗∗ (–1.4) 23.7∗∗ (–3.5) 23.8∗∗ (–3.4) 23.4∗∗ (–3.8) 12.8∗∗∗ (–14.4)
llama-4 15.7∗∗∗ 14.5∗∗∗ (–1.2) 13.0∗∗ (–2.7) 13.1∗∗ (–2.6) 13.8∗∗∗ (–1.9) 7.3∗∗∗ (–8.4)
mistral 5.5∗∗∗ 5.5∗∗∗ (–0.1) 5.7∗∗∗ (+0.1) 4.9∗∗∗ (–0.7) 4.2∗∗∗ (–1.3) 3.9∗∗ (–1.6)
o3 51.5∗∗∗ 52.8∗∗∗ (+1.3) 47.6∗∗ (–3.8) 49.5∗∗∗ (–2.0) 46.8∗∗ (–4.7) 38.6∗∗∗ (–12.9)
o4-mini 41.5∗∗∗ 43.0∗∗∗ (+1.5) 38.3∗∗∗ (–3.2) 38.5∗∗∗ (–3.0) 40.4∗∗∗ (–1.1) 29.1∗∗∗ (–12.4)
qwen3 28.2∗∗∗ 29.3∗∗∗ (+1.1) 27.3∗∗∗ (–0.9) 26.9∗∗∗ (–1.4) 26.5∗∗∗ (–1.7) 14.9∗∗∗ (–13.4)
Note: $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$
Robustness
We evaluated 18 different LLMs on this benchmark, and results are summarized in Table 1. For each variation of the model, we also conducted a two-proportion z-test to measure if the accuracy rate decreases significantly as compared to the original. Statistically significant differences are indicated using standard notation (*$p<0.1$, **$p<0.05$, ***$p<0.01$). We also computed 95 % CI (See Appendix E Figure 4) and robustness metrics $R$ (See Appendix E Figure 7), and all models, especially those performed well on the original set, got a low robustness score.
@@
Transformation-wise Breakdown
We now present a detailed comparison of model robustness across the five variant types: Descriptive, Confusing, Misleading, Garbled String, and Kernel Variant.
Descriptive Long (DL)
@@
This transformation had the smallest impact on accuracy across all models. The average accuracy drop is marginal, and most drops are not statistically significant. Some models such as o3 (+1.3), o4-mini(+1.5), and Qwen3-235B (+1.1) even improved slightly. As expected, changing variables to descriptive names tends to preserve accuracy.
Confusing (DLC)
@@
The addition of semantically meaningless but structurally longer variable names caused a moderate decrease in accuracy. Models like Claude-opus-4 (–4.3**) and GPT-4o-mini (–1.8**) showed statistically significant drops.
Misleading (DLM)
@@
Perturbing the questions by replacing variable names with misleading strings has a strong impact on mathematical reasoning accuracy. Nearly all models experienced a significant drop, with several exceeding –5 percentage points. Notably, Claude-Opus-4 (–4.8**), Gemini-2.5-pro (–5.6***), and Claude-Sonnet-4 (–4.4**) were among the most heavily affected.
Garbled String (GS)
@@
Replacing variable names with random character strings consistently degraded performance across all LLMs. Every model lost accuracy, and more than half of them had drops that reached statistical significance. Models such as Gemini-2.5-pro (–5.4***), Claude-Sonnet-4 (–4.9***), and Gemini-2.5-flash-lite (–4.7***) suffered the largest declines.
Kernel Variant (KV)
@@
Kernel variants—which keep each question’s mathematical structure but replace constants and expressions with different values—led to the sharpest decline overall. All models experienced large drops, often in the range of –5 to –15 points, with Grok4 (–14.6***), Gemini-2.5-flash (–15.2***), and Gemini-2.5-pro (–14.9***) showing the steepest declines.
Overall, these results highlight that even state-of-the-art LLMs struggle to maintain consistent performance under simple, semantics-preserving transformations. These findings suggest that LLMs may not consistently invoke their mathematical reasoning abilities, even when such knowledge is likely embedded in their parameters. Instead, performance appears sensitive to superficial cues such as variable names or textual formatting—features that should be irrelevant to a mathematically robust solution.
Error Taxonomy
@@
Our grading script returns a brief comment for every incorrect answer. Using these comments, we grouped errors into four categories: Symbol Confusion, Step Omission, Arithmetic, and Logic Hallucination. Figure 3 shows that the relative frequency of these error types is nearly identical across variants; logic hallucinations dominate, accounting for roughly three-fifths of all wrong answers regardless of prompt wording. Thus, the accuracy drop is distributed across all categories rather than driven by a single one, confirming that mathematically equivalent perturbation consistently degrades LLM performance.
Figure 3: Error Composition Ratio across Variants
Discussion
Key Findings
Symbol-level perturbations cause substantial drops.
@@ Discussion
Across the four surface variants—DL, DLC, DLM, and GS—merely renaming variables lowers accuracy by $3\text{--}5$ pp on average; for example, Gemini-2.5-pro falls from 78.3% to 72.9% (–5.4 pp; see Table 1). This indicates that today’s SOTA models still rely on lexical “semantic anchors” rather than fully abstract proof structures.
Maintaining structure but resampling parameters is even harsher.
@@
The Kernel Variant (KV) simultaneously resamples all mutable constants while preserving the original reasoning skeleton. Accuracy losses reach $\approx 10$ pp; OpenAI O3 declines from 48.8% to 38.5% (–10.3 pp), showing that grasping a solution pattern does not automatically translate to parameter-invariant reasoning ability.
$R_{\text{global}}$ reveals fine-grained brittleness.
@@
We compute $R_{\text{surf}},R_{\text{core}},R_{\text{global}}$ where $R(\cdot,\cdot)$ is the SD–normalized robustness metric. Because it exponentially penalizes rare but catastrophic flips, $R_{\text{global}}$ tracks effective robustness more faithfully than a plain hard/easy accuracy ratio.
Implications
@@
Evaluation. The GAP framework generates an (in principle) unbounded supply of semantically equivalent test items, limiting future benchmark leakage and mitigating leaderboard inflation.
@@
Training. Results suggest curriculum fine-tuning that explicitly randomizes (i) symbol identities and (ii) numeric parameters, instead of simply enlarging pre-training corpora.
@@
Security. Surface-level fragility implies that production systems can be prompt-injected with mathematically innocuous renamings—highlighting the need to integrate robustness checks into red-team pipelines.
Practical Recommendations
Our study suggests that some strategies such as the following may potentially inprove the performance of LLMs on math reasoning tasks.
  1. 1.
    Data augmentation. Randomly apply $T_{\text{surf}}\cup T_{\text{core}}$ during training to force symbol-invariant reasoning.
  2. 2.
    Symbol binding. Separate identifier tokens from literal tokens (e.g., via a learnable symbol table) inside the Transformer.
  3. 3.
    Hybrid reasoning. Embed SMT/CAS validators into decoding (e.g., value-head alignment) to tighten logical consistency.
Ethical Considerations & Societal Impact
Our benchmark is released under a non-commercial licence with variants and auto-graders only; raw solutions remain withheld. This transparency enables reproducible stress tests while limiting the risk of seeding training corpora with answer keys. Nonetheless, the same techniques could craft adversarial prompts that mislead automated theorem provers, so we encourage multi-agent verification in high-stakes deployments.
Limitations
LLM‑generated variants and verifier bias.
The O3‑based generator and five‑judge verifier bound the chance of accepting a flawed core‑step variant to $\delta<10^{-10}$. Because all judges share the same architecture, rare shared oversights may persist; future versions will mix in symbolic provers or heterogeneously trained LLMs to strengthen the safety net.
Automatic‑grading noise.
Our LLM rubric delivers $\approx 97\%$ precision/recall, leaving a uniform $\lesssim 3\%$ label error that has so far not altered model rankings. Multi‑agent grading or targeted human audits on disagreement cases could tighten this bound.
Residual data leakage.
Kernel resampling and systematic symbol renaming greatly reduce verbatim memorization, yet structurally similar proofs in pre‑training data remain a remote possibility. Hash‑challenge prompts or evaluation‑time watermarking would help detect such leakage.
Metric interpretability.
The SD‑normalized robustness score $R$ is scale‑free and differentiable, but its exponential compression can mask item‑level failure modes. We therefore pair $R$ with distributional diagnostics such as Batched Drop Curves to improve transparency.
Related Work
@@
There have been multiple benchmarks for evaluating the mathematical-reasoning capabilities of large language models (LLMs). Early math‐reasoning benchmarks such as MATH(1.25 k problems), and GSM8K(8.5 k problems), revealed basic arithmetic/algebra skills. (Hendrycks et al. 2021; Cobbe et al. 2021). But their difficulty is now saturated as LLMs scale. For instance, with prompting strategies such as DUP, GPT-4 attains 97.1% accuracy on GSM8K (Zhong et al. 2025). This ceiling at the high-school-competition level motivated the creation of a new generation of harder benchmarks.
@@
Subsequent benchmarks target harder problems. Omni-MATH contributes 4 428 rigorously annotated Olympiad-level problems (Gao et al. 2024). Likewise, OlympiadBench provides a bilingual, multimodal benchmark of 8 476 Olympiad-level math and physics problems with expert step-by-step solutions (He et al. 2024). The cross-disciplinary benchmark ARB consist questions in mathematics, physics, biology, chemistry, and law, with a rubric-based self-grading protocol (Sawada et al. 2023). Some other benchmarks focuses specifically on formal proof. MiniF2F supplies 488 Olympiad-level problems formalized in multiple proof assistants (Zheng, Han, and Polu 2022). Putnambench, offers 1 692 rigorously hand-crafted formalizations of Putnam Competition problems(Tsoukalas et al. 2024).
@@
Nevertheless, recent studies warn that scores on many NLP benchmarks may be artificially inflated by data contamination, when LLMs are trained on the benchmark questions. Sainz et al. point out that many benchmarks may be inflated because large language models often memorize test data seen during pre-training (Sainz et al. 2023). Balloccu et al. conduct a systematic audit of data leakage for closed-source LLMs and estimate that roughly 4.7 million test examples from 263 datasets were likely exposed to the models (Balloccu et al. 2024).
@@
In order to obtain a more robust evaluation of LLMs’ reasoning capabilities, it is important to prevent data leakage. One approach is to create original questions. FrontierMath, for example, addresses this with a rigorously curated benchmark of hundreds of original, expert-level mathematics problems spanning fields from number theory to algebraic geometry (Glazer et al. 2024). Putnam-AXIOM takes this approach and comprises 522 challenging problems from the William Lowell Putnam Competition plus 100 programmatically generated functional variations, furnishing a contamination-resilient benchmark (Gulati et al. 2025).
@@
Another method to deal with data leakage is introducing contrast sets-small, label-changing perturbations of existing test instances—to probe a model’s local decision boundary (Gardner et al. 2020). Huang et al. construct MATH-Perturb that applies simple and hard perturbations to 279 level-5 MATH problems and discover that models suffer 12–16 percentage-point drops on the hard variant(Huang et al. 2025).Shalyt et al. complement this line of work with ASyMOB, a 17 k-problem benchmark whose algebra-focused numerical and symbolic perturbations reveal performance drops of up to 70 percentage points, highlighting models’ fragility under such stress tests (Shalyt, Elimelech, and Kaminer 2025). Similarly, Yu et al. propose Math-RoB, a synthetic benchmark that enables robustness evaluation against the data contamination through an instruction-based approach(Yu et al. 2025). These efforts either focus on specific aspects that limit generalizability, are based on benchmarks that are too easy for current models, or introduce transformations that are not mathematically equivalent, thus confounding true robustness evaluation.
@@
Conclusion & Future Work
In this paper, we introduced the Generalization–and–Perturbation (GAP) framework. By instantiating GAP on all 1 051 Putnam Competition questions we produced the 6 306‑question PutnamGAP benchmark. A zero‑shot evaluation of 18 commercial and open‑source LLMs revealed sharp and consistent accuracy drops. These results expose a clear robustness gap that leaderboard scores on unperturbed datasets have so far not shown.
The findings highlight three actionable directions.

### v2

@@ 5.1 Robustness
5.2 Transformation-wise Breakdown
@@
Descriptive Long (DL) The impact of this transformation is smallest overall: drops are marginal and mostly not significant. Some models such as o3 (+1.3), o4-mini(+1.5), and Qwen3-235B (+1.1) even improved slightly. This indicates that descriptive renaming preserving accuracy.
@@
Confusing (DLC) Long, semantically meaningless variable names moderately reduce accuracy. Models like Claude-opus-4 (–4.3***) and GPT-4o-mini (–1.8***) showed significant drops.
@@
Misleading (DLM) Replacing variables with misleading strings strongly hurts math accuracy. Nearly all models experienced a significant drop. Notably, Claude-Opus-4 (–4.8***), Gemini-2.5-pro (–5.6***), and Claude-Sonnet-4 (–4.4***) were among the most heavily affected.
@@
Garbled String (GS) Random character strings consistently degrade performance: every model loses accuracy, over half significantly. Models such as Gemini-2.5-pro (–5.4***), Claude-Sonnet-4 (–4.9***), and Gemini-2.5-flash-lite (–4.7***) suffered the largest declines.
@@
Kernel Variant (KV) Kernel variants—which keep each question’s mathematical structure but replace constants and expressions with different values—led to the sharpest decline overall. All models experienced large drops, often in the range of –5 to –15 points, with Grok4 (–14.6***), Gemini-2.5-flash (–15.2***), and Gemini-2.5-pro (–14.9***) showing the steepest declines.
Overall, state-of-the-art LLMs show inconsistent performance under semantics-preserving transformations and appear sensitive to superficial cues. This is consistent with the possibility that part of their gains reflects data-leakage–related memorization rather than stable mathematical reasoning. The pattern persists across topics and problem classes: bar plots with $95\%$ CIs (Appendix D, fig. 4) and per-topic/per-class breakdowns (Appendix D, figs. 7-8) show similar robustness gaps across Algebra/Analysis/NT/Combinatorics/Geometry and for both proof and calculation items.
5.3 Error Taxonomy
@@
Our grading script returns a brief comment for every incorrect answer. Using these comments, we grouped errors into four categories: Symbol Confusion, Step Omission, Arithmetic, and Logic Hallucination. Figure 5 in Appendix D shows that the relative frequency of these error types is nearly identical across variants; logic hallucinations dominate, accounting for roughly three-fifths of all wrong answers regardless of prompt wording. Thus, the accuracy drop is distributed across all categories rather than driven by a single one, confirming that mathematically equivalent perturbation consistently degrades LLM performance.
5.4 External Validation
We applied our surface-renaming protocols—DLC and GS—to ALG514  (kus, 2014). Accuracy decreased from Base 93.6% to DLC 90.9% ($\Delta=-2.7$ pp) and GS 89.3% ($\Delta=-4.3$ pp); McNemar tests (Base vs DLC: b=24,c=10,p=0.024b{=}24,c{=}10,p{=}0.024; Base vs GS: b=35,c=13,p=0.002b{=}35,c{=}13,p{=}0.002). These statistically significant drops indicate that GAP’s surface-renaming stress tests generalize to other math datasets and reveal nontrivial sensitivity to variable renaming.
6  Discussion
6.1 Key Findings
The proposed GAP framework allowed us to make the following new findings about the behavior of LLMs in performing mathematical reasoning:
@@ 6.1 Key Findings
Symbol-level perturbations cause substantial drops. Across the four surface variants—DL, DLC, DLM, and GS—merely renaming variables lowers accuracy by $3\text{--}5$ pp on average; for example, Gemini-2.5-pro falls from 78.3% to 72.9% (–5.4 pp; see Table 1). This indicates that today’s SOTA models still rely on lexical “semantic anchors” rather than fully abstract proof structures.
@@
Maintaining structure but resampling parameters is even harsher. The Kernel Variant (KV) simultaneously resamples all mutable constants while preserving the original reasoning skeleton. Accuracy losses reach $\approx 10$ pp; OpenAI O3 declines from 48.8% to 38.5% (–10.3 pp), showing that grasping a solution pattern does not automatically translate to parameter-invariant reasoning ability.
@@
$R_{\text{global}}$ reveals fine-grained brittleness. We compute $R_{\text{surf}},R_{\text{para}},R_{\text{global}}$ where $R(\cdot,\cdot)$ is the SD–normalized robustness metric. Because it exponentially penalizes rare but catastrophic flips, $R_{\text{global}}$ tracks effective robustness more faithfully than a plain hard/easy accuracy ratio.
Takeaway. Across capable models we consistently observe $R_{\text{para}}<R_{\text{surf}}$, and we summarize stress-type invariance via $R_{\text{global}}=\sqrt{R_{\text{surf}}R_{\text{para}}}$; interpreting 1R1-R as penalty mass highlights non-trivial fragility even when raw accuracy is high.
6.2 Implications
@@
A novel evaluation methodology: The GAP framework provides a novel methodology for analyzing and evaluating the robustness of LLMs’ reasoning capacity by generating an (in principle) unbounded supply of semantically equivalent test items, which can limit future benchmark leakage and mitigate leaderboard inflation.
@@
Improving robustness via curriculum fine-tuning: Our results suggest curriculum fine-tuning that explicitly randomizes (i) symbol identities and (ii) numeric parameters, instead of simply enlarging pre-training corpora. That is, we can leverage the GAP framework to augment data for fine-tuning a model to improve robustness.
@@
Detecting potential security concerns: Surface-level fragility implies that production systems can be prompt-injected with mathematically innocuous renamings—highlighting the need to integrate robustness checks into red-team pipelines. Our evaluation framework enables such risk analysis before deploying any production system.
Reporting. We recommend reporting bootstrap CIs for $R_{b}$ together with per-item histograms of SD-normalized drops $d_{j}=(e_{j}-h_{j})/\sigma$; these visualize tail-risk (rare catastrophic flips) that raw accuracy masks and make robustness audits reproducible.
7  Related Work
@@
There have been multiple benchmarks for evaluating the mathematical-reasoning capabilities of large language models (LLMs). Early math‐reasoning benchmarks such as MATH(1.25 k problems) (Hendrycks et al., 2021), and GSM8K(8.5 k problems) (Cobbe et al., 2021), revealed basic arithmetic/algebra skills. But their difficulty is now saturated as LLMs scale. For instance, with prompting strategies such as DUP, GPT-4 attains 97.1% accuracy on GSM8K (Zhong et al., 2025). This ceiling at the high-school-competition level motivated the creation of a new generation of harder benchmarks.
@@
Subsequent benchmarks target harder problems. Omni-MATH contributes 4 428 rigorously annotated Olympiad-level problems (Gao et al., 2024). Likewise, OlympiadBench provides a bilingual, multimodal benchmark of 8 476 Olympiad-level math and physics problems with expert step-by-step solutions (He et al., 2024). The cross-disciplinary benchmark ARB consist questions in mathematics, physics, biology, chemistry, and law, with a rubric-based self-grading protocol (Sawada et al., 2023). Some other benchmarks focuses specifically on formal proof. MiniF2F supplies 488 Olympiad-level problems formalized in multiple proof assistants (Zheng et al., 2022). Putnambench, offers 1 692 rigorously hand-crafted formalizations of Putnam Competition problems (Tsoukalas et al., 2024).
@@
Nevertheless, recent studies warn that scores on many NLP benchmarks may be artificially inflated by data contamination, when LLMs are trained on the benchmark questions. Sainz et al. (2023) point out that many benchmarks may be inflated because large language models often memorize test data seen during pre-training.  Balloccu et al. (2024) conduct a systematic audit of data leakage for closed-source LLMs and estimate that roughly 4.7 million test examples from 263 datasets were likely exposed to the models.
@@
In order to obtain a more robust evaluation of LLMs’ reasoning capabilities, it is important to prevent data leakage. One approach is to create original questions. FrontierMath, for example, addresses this with a rigorously curated benchmark of hundreds of original, expert-level mathematics problems spanning fields from number theory to algebraic geometry (Glazer et al., 2024). Putnam-AXIOM takes this approach and comprises 522 challenging problems from the William Lowell Putnam Competition plus 100 programmatically generated functional variations, furnishing a contamination-resilient benchmark (Gulati et al., 2025).
@@
Another method to deal with data leakage is introducing contrast sets-small, label-changing perturbations of existing test instances—to probe a model’s local decision boundary (Gardner et al., 2020). Huang et al. (2025) construct MATH-Perturb that applies simple and hard perturbations to 279 level-5 MATH problems and discover that models suffer 12–16 percentage-point drops on the hard variant. Shalyt et al. (2025) complement this line of work with ASyMOB, a 17k-problem benchmark whose algebra-focused numerical and symbolic perturbations reveal performance drops of up to 70 percentage points, highlighting models’ fragility under such stress tests. Similarly, Yu et al. (2025) propose Math-RoB, a synthetic benchmark that enables robustness evaluation against the data contamination through an instruction-based approach. These efforts either focus on specific aspects that limit generalizability, are based on benchmarks that are too easy for current models, or introduce transformations that are not mathematically equivalent, thus confounding true robustness evaluation.
@@
8  Conclusion & Future Work
Robust reasoning is required in many applications of LLMs. In this paper, we proposed a novel Generalization–and–Perturbation (GAP) framework for analyzing and evaluating robustness of LLMs’ reasoning capacity. By instantiating GAP on all 1,051 Putnam Competition questions we produced the 6,306‑question PutnamGAP benchmark. A zero‑shot evaluation of 18 commercial and open‑source LLMs revealed sharp and consistent accuracy drops. These results expose a clear robustness gap that leaderboard scores on unperturbed datasets have so far not shown.
Our findings highlight three actionable directions.